# Synthesizing data products, mathematical models, and observations for lake temperature forecasting

### v2

@@ 3 GP surrogates for forecasting
There are two challenges to extending the proof-of-concept in Section 2. The first is conceptual: pivoting the modeling framework toward forecasts. We wish to synthesize computer model runs driven by an ensemble of “plausible weather futures” so that they may be compared to what actually happened, as observed in the field. We need a model for how GLM – driven by NOAA ensembles over extended horizons [Figure 3] – relates to temperatures observed by sensors at FCR. The second challenge is technological, centering around modeling fidelity and scale. Thirty-day horizons, and 31 NOAA ensemble members over the course of multiple years (and depths) where observational data has been obtained, represent a very large, and heteroskedastic [Figure 3], simulation campaign. Both, large scale GP modeling (Heaton et al., 2019) and accommodating input-dependent variance for stochastic simulation (Baker et al., 2022), are on the frontier of the surrogate modeling landscape. Here we describe a setup that provides enough flexibility for our NOAA-GLM at FCR while remaining practical, off-loading work to modern libraries as much as possible, and computationally tractable.
@@ 3.1 Modeling a forecasting apparatus
This can be repeated for each horizon $h$ of interest. Such a setup has much to recommend it. It is modular, in that it uses off-the-shelf tools, and it is computationally tractable because each GP is fit to moderately sized training data, keeping the matrix decompositions manageable. But such modeling is inefficient statistically because separate, independent fits do not account for correlations across horizon, which are clearly evident in Figure 3, both in mean and in variance. Furthermore, UQ will be less accurate, because through averaging the 31 ensemble GLM simulations, variance information is lost. Finally, since $t$ represents Julian day, adjustments are required to accommodate multiple years of data.
Toward a more holistic model, we propose the following. Let $\mathbf{X}_{\_}{n_{\_}M}$ represent the matrix of inputs as in Section 2, but now augmented with a horizon column $h\in\{1,\dots,h_{\_}{\max}=30\}$, an ensemble member column $\xi\in\{1,\dots,31\}$, and a year column $\lambda$. The $t$ column would be interpreted as the day the forecast was made; the last day for which observational data (i.e., the truth) is known, i.e., $t-h$ from the description above, re-labeled. Columns $\xi$ and $\lambda$ are merely included to index the data, and we shall provide more detail about how these affect modeling momentarily. We may create such a matrix, i.e., beyond the time and depth columns previously contained in $\mathbf{X}_{\_}{n_{\_}M}$ from Section 2, via Cartesian product so that $n_{\_}M\leftarrow n_{\_}M\times 30\times 31\times 3$. Now let $D_{\_}{n_{\_}M}=(\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M})$ collect simulation data relating time, horizon, depth, and ensemble member to GLM-simulated temperature(s). Supposing we could surrogate model these simulations, we could again follow the recipe in Section 2, filtering surrogate fits and bias corrections through to forecasts. But there are logistical challenges with $n_{\_}M$ in the millions and with heteroskedastic NOAA-GLM runs [Figure 3].
@@ 3.2 Heteroscedastic surrogate modeling
We begin by discarding the ensemble column $\xi$ in $\mathbf{X}_{\_}{n_{\_}M}$ and treating forecasts from each $\xi$ as independent conditional on time and depth. This modeling assertion is purely for computational convenience; it is not faithful to the simulation mechanism. While NOAA ensemble members are independent of one another, each realization evolves in a time-dependent manner. Forecasts from the prior day or depth for a specific ensemble member might contain valuable data for predicting the following day or depth for that same member, and potentially others. Nonetheless, independence aligns with the essence of ensemble modeling (Thomas et al., 2020).
This “modeling hack” means that each of 31 ensemble forecasts may be treated as replicates. Baker et al. (2022) explain that replicates bring two kinds of efficiency. One is computational: the number of sufficient statistics are fewer – 31-fold fewer in this instance. The other is statistical: replicates allow a pure glimpse at variance, without needing to first separate signal from noise. This is especially important when noise levels are changing along with the mean, in a potentially nonlinear way as in Figure 3. The most effective methods for modeling such nonlinear heteroskedasticity (e.g., Goldberg, Williams and Bishop, 1997; Ankenman, Nelson and Staum, 2008; Quadrianto et al., 2009; Lázaro-Gredilla and Titsias, 2011; Binois, Gramacy and Ludkovski, 2018) leverage replication. Stochastic kriging (SK; Ankenman, Nelson and Staum, 2008), one of the first of these, is ideal for our setting, in part because it cuts corners relative to other methods [cited above]; e.g., (a) being moment-based (rather than likelihood-based); and (b) not jointly modeling the mean and variance process(es). Indeed, both are valid concerns when the degree of replication is low. Ankenman, Nelson and Staum (2008) recommend at least ten replicates for each otherwise unique input. We have 31, more than plenty on that account. Perhaps more importantly, SK is uniquely situated to interface with our second layer of approximation in Section 3.3.
@@
Finally, store these in $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, respectively, to line up with the rows of $\bar{\mathbf{X}}_{\_}n$. Note that we are dropping the $M$ subscripts with the understanding that we are only using this on the computer model runs. If the field data also had replicates we could additionally make use of them there, but our sensors record only one measurement per day.
@@
Now consider new/forecasting coordinates $\mathcal{X}$ that may be of interest, like for a particular day $t$ and multiple horizons $h$ into the future, or representing hindcasts to compare to to sensor measurements discussed further in Section 3.4. Let $s(\mathcal{X})=\mathrm{Diag}(T^{-1}(\mu^{(v)}(\mathcal{X})))$ capture these back-transformed variance predictions – they are the predictive analog of Eq. (8). Using these, via $\mathbf{S}_{\_}n$ and $\mathcal{GP}(D^{(m)})$ where $D^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$, the SK equations are
@@
Figure 5: 2d slices of predictive mean $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ (left) and variance $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$ (middle) from the (heteroscedastic) GLM surrogate over depth (m) and horizon (days). Together, these summarize a hindcast for 2022-02-19. The right panel shows a 1m-depth slice with GLM ensemble forecasts (light gray lines), SK surrogate (darker, solid line) 90% CIs and PI (dotted and dashed lines). Observations are overlaid as darker colored dots.
To illustrate, consider Figure 5 which is derived from fits of two GPs, for mean and variance, to a simulation campaign spanning three years over a thirty-day horizon. The data represented in the plot comprises a 2d slice of predictions from those two fits corresponding to a forecast beginning on 2022-02-19, offering a surrogate for the NOAA-GLM hindcast beginning on that date. Focusing first on the left panel, showing $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ from Eq. (9), observe that temperature is decreasing with increasing depth, and also generally increasing over horizon, most prominently at the surface (low depths). Both relationships are clearly nonlinear. Now looking at the middle panel providing $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}=\mathrm{Diag}(\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X}))$ we see a similar relationship in the spread of possible temperatures. Variance is much lower at lower depths, even at longer horizons, because these depths are more insulated from weather at the surface. The right panel of the figure shows a 1d slice of the surface(s) 1m depth along with additional information. Gray lines indicate raw NOAA-GLM simulations (i.e., a subset of $D_{\_}{n_{\_}M}$). Filled purple dots indicate sensor observations. These latter values are not part of the analysis at this time and are provided for reference only. Observe that the dashed orange 90% PIs, derived from $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ and $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$, have appropriate coverage. Dotted lines indicate 90% CIs on the mean that are derived by eliminating the $s(\mathcal{X})$ term in Eq. (10), as an analogue to Eq. (5). It is this squared standard error which is crucial to our bias-correcting analysis described in Section 3.4. Finally, observe that NOAA-GLM has a cold bias compared to the field data.
@@ 3.3 GP surrogates and scalability
The surrogates behind the views in Figures 5 were constructed using the approximate methods we describe here. Scalable GP approximation has been an active field of late. Techniques include compactly supported kernels (Gneiting, 2002; Kaufman et al., 2011), fixed-rank kriging (Cressie and Johannesson, 2008) inducing points (Banerjee et al., 2008; Quinonero-Candela and Rasmussen, 2005; Hensman, Fusi and Lawrence, 2013) local divide-and-conquer (Emery, 2009; Gramacy, 2016) and the Vecchia/nearest-neighbor approximation (e.g., Datta et al., 2016; Katzfuss et al., 2020a) The so-called scaled Vecchia approximation (SVecchia; Katzfuss, Guinness and Lawrence, 2022) seems particularly well-suited to our setting because it allows coordinate-wise lengthscale estimation and scales in flops that are quasi-linear in $n$. Crucially, an open-source implementation is provided for R: https://github.com/katzfuss-group/scaledVecchia via GpGp (Guinness, 2018) and GPVecchia (Katzfuss et al., 2020b) packages. We provide a review of the Vecchia approximation, and some implementation details in Appendix A.1.
@@
Vecchia has never been used, and nor to our knowledge has any other large-scale GP approximation, to model input-dependent variances, like through SK (Section 3.2). Fitting second moments $\mathcal{GP}(D_{\_}n^{(v)}$) in order to get a predictor for $T(\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathcal{X}})\})$ is straightforward, because this amounts to feeding in (transformed) $s_{\_}i$-values (Eq. 7) as responses along with unique inputs $\bar{\mathbf{x}}_{\_}i$. The visual in the right panel of Figure 5 shows $\mu^{(v)}_{\_}n(\mathcal{X})$. Using Vecchia to fit the mean GP conditional on those variances (left panel of Figure 5) is harder because the R library we use does not accommodate a vector of nuggets, which comprises $\mathbf{S}_{\_}n$ in Eqs. (910). So we ended up implementing the following workaround instead.
@@
We begin by fitting first moments with an otherwise ordinary Vecchia approximated GP, i.e., $\mathcal{GP}(D_{\_}n^{(m)})$, where $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$. Let $\mu^{(m)}_{\_}n(\mathcal{X})$ denote the predictive mean of $\mathcal{GP}(D_{\_}n^{(m)})$. Note that $\mu^{(m)}_{\_}n\neq\mu^{\mathrm{SK}}_{\_}n$ from Eq. (9) because the former utilizes a covariance structure with a single nugget, i.e., $\tau^{2}(k(\bar{\mathbf{X}}_{\_}n)+g\mathbb{I}_{\_}n)$ following Eq. (4), and the latter uses $\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n$. However, they are equal in expectation because GPs furnish unbiased predictors (see, e.g., Santner et al., 2018). Nevertheless, the two can be quite different if the degree of replication is high because of the $\mathbf{n}$-vector in the denominator of Eq. (8). Specifically, if the replication degree varies significantly from one location to another, for instance $n_{\_}i\gg n_{\_}j$ for $\bar{\mathbf{x}}_{\_}i$ and $\bar{\mathbf{x}}_{\_}j$, then the predictions from $\mu^{\mathrm{SK}}_{\_}n(\mathbf{x})$ for $\mathbf{x}$ in the vicinity of $\bar{\mathbf{x}}_{\_}j$, the site with fewer replicates, could be considerably more precise than $\mu^{(m)}_{\_}n(\mathbf{x})$, which does not know to place less trust in $\bar{y}_{\_}j$ than in other locations, such as near $\bar{\mathbf{x}}_{\_}i$. However, in our setting, we have uniform replication, with all $n_{\_}i=31$. In a study described in Appendix A.3, and summarized in Figure 13, we find there is essentially no difference between $\mu^{(m)}_{\_}n$, via Vecchia, and $\mu^{\mathrm{SK}}_{\_}n$.
@@
Approximating what SK provides for the variance of the first moment, $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ in Eq. (10), with Vecchia, follows a similar thread. Whereas our development of the mean leverages the unbiased nature of Gaussian expectations, second moments follow standard calculations of Gaussian variances. Crucially, we do not need the entire covariance structure $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ for bias correction in Section 3.4; we need only the diagonal. Those “variances of the mean” are Gaussian standard errors, i.e., $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}/n$, again leveraging uniform replication: $n_{\_}i=31$. Our Vecchia fit to second moments provides an estimate of $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ so all we need to do is divide this quantity by $n_{\_}i$, for any $i$, to get an estimate of the variance of the mean. However, we prefer a more conservative application of this in an SK-approximating context to avoid under-covering out-of-sample. We found that using the 95th quantile of the predictive mean obtained from $\mathcal{GP}(D_{\_}n^{(v)})$, notated as $\mu^{(v)95}_{\_}n(\mathbf{x})$, added to the variance of the first moments’ fit ($\sigma^{2m}_{\_}n(\mathbf{x})$) provided sufficient coverage. That is, we use
@@
Beyond missing off-diagonal covariance terms, this is an inefficient estimator compared to the ideal SK quantity in Eq. (10) because it does not borrow strength nearby in the input space. (This is what the term $-\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\mathbf{X}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})$ in that expression is doing.) So in $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ we are getting over-inflated, conservative UQ because the Veccia software we are using does not provide a full covariance.
Appendix A.2A.3 provides two examples showing that this approach achieves the desired results in a benchmark/toy example and on GLM runs, respectively. These are intended to augment other visuals in the paper, which are all derived from Vechhia approximations.

### v3

@@ 3 GP surrogates for forecasting
There are two challenges to extending the proof-of-concept in Section 2. The first is conceptual: pivoting the modeling framework toward forecasts. We wish to synthesize computer model runs driven by an ensemble of “plausible weather futures” so that they may be compared to what actually happened, as observed in the field. We need a model for how GLM – driven by NOAA ensembles over extended horizons [Figure 3] – relates to temperatures observed by sensors at FCR. The second challenge is technological, centering around modeling fidelity and scale. Thirty-day horizons, and 31 NOAA ensemble members over the course of multiple years (and depths) where observational data has been obtained, represent a very large, and heteroskedastic [Figure 3], simulation campaign. Both large-scale GP modeling (Heaton et al., 2019) and accommodating input-dependent variance for stochastic simulation (Baker et al., 2022) are on the frontier of the surrogate modeling landscape. Here we describe a setup that provides enough flexibility for our NOAA-GLM at FCR while remaining practical, off-loading work to modern libraries as much as possible, and computationally tractable.
@@ 3.1 Modeling a forecasting apparatus
This can be repeated for each horizon $h$ of interest. Such a setup has much to recommend it. It is modular, in that it uses off-the-shelf tools, and it is computationally tractable because each GP is fit to moderately sized training data, keeping the matrix decompositions manageable. Such modeling is inefficient statistically because separate, independent fits do not account for correlations across horizon, which are clearly evident in Figure 3, both in mean and in variance. Furthermore, UQ will be less accurate, because through averaging the 31 ensemble GLM simulations, variance information is lost. Finally, since $t$ represents Julian day, adjustments are required to accommodate multiple years of data.
Toward a more holistic model, we propose the following. Let $\mathbf{X}_{\_}{n_{\_}M}$ represent the matrix of inputs as in Section 2, but now augmented with a horizon column $h\in\{1,\dots,h_{\_}{\max}=30\}$, an ensemble member column $\xi\in\{1,\dots,31\}$, and a year column $\lambda$. The $t$ column would be interpreted as the day the forecast was made; the last day for which observational data (i.e., the truth) is known, i.e., $t-h$ from the description above, re-labeled. Columns $\xi$ and $\lambda$ are merely included to index the data, and we shall provide more detail about how these affect modeling momentarily. We may create such a matrix, i.e., beyond the time and depth columns previously contained in $\mathbf{X}_{\_}{n_{\_}M}$ from Section 2, via Cartesian product so that $n_{\_}M\leftarrow n_{\_}M\times 30\times 31\times 3$. Now let $D_{\_}{n_{\_}M}=(\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M})$ collect simulation data relating horizon, ensemble member, time, and depth to GLM-simulated temperature(s). Supposing we could surrogate model these simulations, we could again follow the recipe in Section 2, filtering surrogate fits and bias corrections through to forecasts. There are still logistical challenges with $n_{\_}M$ in the millions and with heteroskedastic NOAA-GLM runs [Figure 3].
@@ 3.2 Heteroscedastic surrogate modeling
We begin by discarding the ensemble column $\xi$ in $\mathbf{X}_{\_}{n_{\_}M}$ and treating forecasts from each $\xi$ as independent conditional on time and depth. This modeling assertion is purely for computational convenience; it is not faithful to the simulation mechanism. While NOAA ensemble members are independent of one another, each realization evolves in a time-dependent manner. Nonetheless, independence aligns with the essence of ensemble modeling (Thomas et al., 2020). The 31-member NOAA ensemble allows for the measurement of a variety of environmental conditions, whose variability grows over time. We apply these to GLM simulations, as depicted in Figure 3.
This “modeling hack” means that each of 31 ensemble forecasts may be treated as replicates. Baker et al. (2022) explain that replicates bring two kinds of efficiency. One is computational: the number of sufficient statistics are fewer – 31-fold fewer in this instance. The other is statistical: replicates allow a pure glimpse at variance, without needing to first separate signal from noise. This is especially important when noise levels are changing along with the mean, in a potentially nonlinear way as in Figure 3. The most effective methods for modeling such nonlinear heteroskedasticity (see, e.g., Binois, Gramacy and Ludkovski, 2018) leverage replication. Stochastic kriging (SK; Ankenman, Nelson and Staum, 2008), one of the first of these, is ideal for our setting, in part because it cuts corners relative to other methods [cited above]; e.g., (a) being moment-based (rather than likelihood-based); and (b) not jointly modeling the mean and variance process(es). Indeed, both are valid concerns when the degree of replication is low. Ankenman, Nelson and Staum (2008) recommend at least ten replicates for each otherwise unique input. We have 31.Perhaps more importantly, SK is uniquely situated to interface with our second layer of approximation in Section 3.3.
@@
Finally, store these in $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, respectively, to line up with the rows of $\bar{\mathbf{X}}_{\_}n$. Note that we are dropping the $M$ subscripts with the understanding that we are only using this on the computer model runs.
@@
Now consider new/forecasting coordinates $\mathcal{X}$ that may be of interest, like for a particular day $t$ and multiple horizons $h$ into the future, or representing hindcasts to compare to sensor measurements discussed further in Section 3.4. Let $s(\mathcal{X})=\mathrm{Diag}(T^{-1}(\mu^{(v)}(\mathcal{X})))$ capture these back-transformed variance predictions – they are the predictive analog of Eq. (8). Using these, via $\mathbf{S}_{\_}n$ and $\mathcal{GP}(D^{(m)})$ where $D^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$, the SK equations are
@@
Figure 5: 2d slices of predictive mean $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ (left) and variance $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$ (middle) from the (heteroscedastic) GLM surrogate over depth (m) and horizon (days). Together, these summarize a hindcast for 2022-02-19. The right panel shows a 1m-depth slice with GLM ensemble forecasts (light gray lines), SK surrogate (darker, solid line) 90% CIs and PI (dotted and dashed lines). Observations are overlaid as darker dots.
To illustrate, consider Figure 5 which is derived from fits of two GPs, for mean and variance, to a simulation campaign spanning three years over a thirty-day horizon. The data represented in the plot comprises a 2d slice of predictions from those two fits corresponding to a forecast beginning on 2022-02-19, offering a surrogate for the NOAA-GLM hindcast beginning on that date. Focusing first on the left panel, showing $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ from Eq. (9), observe that temperature is decreasing with increasing depth, and generally increasing over horizon, most prominently at the surface (low depths). Now looking at the middle panel providing $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}=\mathrm{Diag}(\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X}))$ we see a similar relationship in the spread of possible temperatures. Variance is much lower at lower depths, even at longer horizons, because these depths are more insulated from weather at the surface. The right panel of the figure shows a 1d slice at 1m depth. Gray lines indicate raw NOAA-GLM simulations (i.e., a subset of $D_{\_}{n_{\_}M}$). Solid dots indicate sensor observations. Observe that the 90% PIs (dashed lines), derived from $\mu^{\mathrm{SK}}_{\_}n(\mathcal{X})$ and $\sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})^{2}$, have appropriate coverage. Dotted lines indicate 90% CIs on the mean that are derived by eliminating the $s(\mathcal{X})$ term in Eq. (10), as an analogue to Eq. (5). It is this squared standard error which is crucial to our bias-correcting analysis described in Section 3.4. Finally, observe that NOAA-GLM has a cold bias compared to the field data.
@@ 3.3 GP surrogates and scalability
The surrogates behind Figure 5 were actually constructed via GP approximation, which has been an active field of late. The so-called scaled Vecchia approximation (SVecchia; Katzfuss, Guinness and Lawrence, 2022) seems particularly well-suited to our setting because it allows coordinate-wise lengthscale estimation and scales in flops that are quasi-linear in $n$. Crucially, an open-source implementation is provided for R: https://github.com/katzfuss-group/scaledVecchia via GpGp (Guinness, 2018) and GPVecchia (Katzfuss et al., 2020) packages. We provide a review of the Vecchia approximation, and some implementation details in Section A.1 of the Supplementary Material (Holthuijzen et al., 2025).
@@
Vecchia has never been used, and nor to our knowledge has any other large-scale GP approximation, to model input-dependent variances, like through SK (Section 3.2). Fitting second moments $\mathcal{GP}(D_{\_}n^{(v)}$) in order to get a predictor for $T(\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathcal{X}})\})$ is straightforward, because this amounts to feeding in (transformed) $s_{\_}i$-values (Eq. 7) as responses along with unique inputs $\bar{\mathbf{x}}_{\_}i$. The visual in the right panel of Figure 5 shows $\mu^{(v)}_{\_}n(\mathcal{X})$. Using Vecchia to fit the mean GP conditional on those variances (left panel of Figure 5) is harder because the R library we use does not accommodate a vector of nuggets, which comprises $\mathbf{S}_{\_}n$ in Eqs. (910). We implemented the following workaround instead.
@@
We begin by fitting first moments with an otherwise ordinary Vecchia approximated GP, i.e., $\mathcal{GP}(D_{\_}n^{(m)})$, where $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$. Let $\mu^{(m)}_{\_}n(\mathcal{X})$ denote the predictive mean of $\mathcal{GP}(D_{\_}n^{(m)})$. Note that $\mu^{(m)}_{\_}n\neq\mu^{\mathrm{SK}}_{\_}n$ from Eq. (9) because the former utilizes a covariance structure with a single nugget, i.e., $\tau^{2}(k(\bar{\mathbf{X}}_{\_}n)+g\mathbb{I}_{\_}n)$ following Eq. (4), and the latter uses $\tau^{2}k(\bar{\mathbf{X}}_{\_}n)+\mathbf{S}_{\_}n$. However, they are equal in expectation because GPs furnish unbiased predictors (see, e.g., Santner et al., 2018). Nevertheless, the two can be quite different if the degree of replication is high because of the $\mathbf{n}$-vector in the denominator of Eq. (8). Specifically, if the replication degree varies significantly from one location to another, for instance $n_{\_}i\gg n_{\_}j$ for $\bar{\mathbf{x}}_{\_}i$ and $\bar{\mathbf{x}}_{\_}j$, then the predictions from $\mu^{\mathrm{SK}}_{\_}n(\mathbf{x})$ for $\mathbf{x}$ in the vicinity of $\bar{\mathbf{x}}_{\_}j$, the site with fewer replicates, could be considerably more precise than $\mu^{(m)}_{\_}n(\mathbf{x})$, which does not know to place less trust in $\bar{y}_{\_}j$ than in other locations, such as near $\bar{\mathbf{x}}_{\_}i$. However, in our setting, we have uniform replication, with all $n_{\_}i=31$. In a study described in Section A.3 and summarized in Figure A1 of the Supplementary Material (Holthuijzen et al., 2025), we find there is essentially no difference between $\mu^{(m)}_{\_}n$, via Vecchia, and $\mu^{\mathrm{SK}}_{\_}n$.
@@
Approximating what SK provides for the variance of the first moment, $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ in Eq. (10), with Vecchia, follows a similar thread. Whereas our development of the mean leverages the unbiased nature of Gaussian expectations, second moments follow standard calculations of Gaussian variances. Crucially, we do not need the entire covariance structure $\Sigma^{\mathrm{SK}}_{\_}n(\mathcal{X})$ for bias correction in Section 3.4; we need only the diagonal. Those “variances of the mean” are Gaussian standard errors, i.e., $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}/n$, again leveraging uniform replication: $n_{\_}i=31$. Our Vecchia fit to second moments provides an estimate of $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ so all we need to do is divide this quantity by $n_{\_}i$, for any $i$, to get an estimate of the variance of the mean. Importantly, we prefer a more conservative application of this in an SK-approximating context to avoid under-covering out-of-sample. We found that using the 95th quantile of the predictive mean obtained from $\mathcal{GP}(D_{\_}n^{(v)})$, notated as $\mu^{(v)95}_{\_}n(\mathbf{x})$, added to the variance of the first moments’ fit ($\sigma^{2m}_{\_}n(\mathbf{x})$) provided sufficient coverage. That is, we use
@@
Beyond missing off-diagonal covariance terms, this is an inefficient estimator compared to the ideal SK quantity in Eq. (10) because it does not borrow strength nearby in the input space. (This is what the term $-\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\mathbf{X}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})$ in that expression is doing.) So in $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ we are getting over-inflated, conservative UQ because the Vecchia software we are using does not provide a full covariance. Note that $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ from Eq. (11) will give rise to CIs; however, if division by $n_{\_}i$ is omitted, one may obtain PIs. The construction of PIs in this manner is only valid for the GLM surrogate without bias-correction, as will be explained in the next section.
Sections A.2–A.3 of the Supplementary Material (Holthuijzen et al., 2025) provide two examples showing that this approach achieves the desired results in a benchmark/toy example and on GLM runs, respectively. These are intended to augment other visuals in the paper, which are all derived from Vecchia approximations.